# On the Three Demons in Causality in Finance: Time Resolution, Nonstationarity, and Latent Factors

### 28 Dec 2023 v1

@@ 1 Introduction
Another focal point in causal discovery and finance is the issue of data nonstationarity, influencing the outcomes of causal discovery methods applied to financial data. This nonstationarity implies that causal relations and their strengths vary with time (Zhang et al., 2015). Within finance, this phenomenon is commonly referred to as concept drift, signifying the changing distribution of data over time (Lu et al., 2018). Traditional causal discovery methods such as PC (Spirtes et al., 2000), GES (Chickering, 2002), and LiNGAM (Shimizu et al., 2006) were originally designed for the i.i.d. case and did not account for distribution shifts. A recent method, PC-MCI (Runge et al., 2019), is specifically tailored for time series data; however, it still exhibits limitations when confronted with changing causal relations over time.
@@
Our contribution can be summarised as below:
  • We delineate three pivotal issues concerning causality in financial data and undertake a comprehensive exploration on these challenges in a systematic manner.
  • We present well-defined methods tailored to address each of the three identified issues. Our preliminary findings reveal compelling and novel insights, providing a foundation for further exploration.
  • Our empirical study validates the effectiveness of our proposed methods in real-life scenarios in finite sample cases.
@@ 1.1 Concepts and Terminology in Causal Discovery
Structural Causal Model. The Structural Causal Model (SCM) comprises a set of endogenous variables ($V$) and a set of exogenous variables ($U$), interconnected by a set of functions ($F$) that illustrate how the changes of exogenous variables $U$ impact the values of endogenous variables $V$ in the data generation process (Pearl, 2009). For instance, for $x_{i}\in V$, $u_{i}\in U$, and $f_{i}\in F$, the model $x_{i}:=f_{i}(Pa(x_{i}),u_{i}),i=1,\dots,d$, signifies the assignment of the value $x_{i}$ to a function $f_{i}$ of its structural parents $Pa(x_{i})$ and exogenous variable $u_{i}$. For each SCM, a causal graph $\mathcal{G}$ is correspondingly generated by representing each $x_{i}$ with each vertex and adding directing edges from each parent variable in $Pa(x_{i})$ to the child $x_{i}$.
Causal Sufficiency. A set of observed variables is deemed causally sufficient if all common causes of these variables are also observed. In other words, there is no latent confounder in the causal graph and all causally related variables are observed.
Causal Markov Condition. In the causal graph, each variable is independent of its non-descendants conditional on all of its direct causes (parents). This relationship establishes a connection from the causal graph to the statistical properties.
Markov Equivalence Class. A set of causal graphs belong to the same Markov Equivalent Class if they entail the same conditional independence relations among the observed variables, though the specific causal structures may be different.
Faithfulness. A probability distribution P is faithful to a causal graph $\mathcal{G}$ if all conditional independence relations in the probability distribution P are entailed by the causal graph $\mathcal{G}$. This relationship establishes a connection from the statistical properties to the causal graph.
@@ 2 Time Resolution and Aggregation
$\mathbf{X}_{t}=\mathrm{A}\mathbf{X}_{t-1}+\mathbf{e}_{t}$ (1)
where $\mathbf{X}_{t}=(X_{t,1},X_{t,2},\dots,X_{t,n})^{T}$ represents the original data vector, $\mathbf{e}_{t}=(e_{t,1},e_{t,2},\dots,e_{t,n})^{T}$ is the temporally and contemporaneously independent noise vector, and $\mathrm{A}$ is the causal transition matrix.
Furthermore, the temporally aggregated data (observed data)
@@
Merging (1) into (2), we have:
$\tilde{\mathbf{X}}_{t}$ $=\frac{1}{k}\sum_{i=1}^{k}\mathbf{X}_{i+(t-1)k}$ (3)
$=\frac{1}{k}\sum_{i=1}^{k}\left(\mathrm{A}\mathbf{X}_{i+(t-1)k-1}+\mathbf{e}_{i+(t-1)k}\right)$
$=\mathrm{A}\left[\frac{1}{k}\sum_{i=1}^{k}\mathbf{X}_{i+(t-1)k}-\frac{1}{k}\left(\mathbf{X}_{tk}-\mathbf{X}_{(t-1)k}\right)\right]+\frac{1}{k}\sum_{i=1}^{k}\mathbf{e}_{i+(t-1)k}$
@@
$\tilde{\mathbf{X}}_{t}=\mathrm{A}\tilde{\mathbf{X}}_{t}+\tilde{\mathbf{e}}_{t}$ (4)
where $\tilde{\mathbf{e}}_{t}=\frac{1}{k}\sum_{i=1}^{k}\mathbf{e}_{i+(t-1)k}$ is the mixture of independent component vectors $\mathbf{e}_{t}$, and elements in $\tilde{\mathbf{e}}_{t}$ are contemporaneous independent to each other. Clearly, $\tilde{\mathbf{X}}_{t}$ exhibits behaviors resembling a linear instantaneous causal process in the i.i.d. case, and hence instantaneous causal discovery methods can be directly applied to aggregated data to discover meaningful causal relationships.
@@ 3.1 From the Financial Perspective
Figure 1: Scatter Plot of PFE / BA Daily Stock Prices
Figure 2: PFE / BA Daily Stock Prices
@@
Figures 1 and 2 demonstrate the evolving correlation between the stock price time series of Pfizer Inc. (PFE) and The Boeing Company (BA) from 2008 to 2023. This relationship has experienced significant shifts, particularly during the pre- and post-COVID periods. Between 2008 and 2019, a time of economic expansion, both PFE and BA exhibited an upward trend in stock prices. The correlation observed in this era, represented by the darkest green scatters in Figure 1, could be influenced by broader economic factors (confounders). The onset of COVID-19 in early 2020 saw a sharp decline in BA’s stock price. However, during 2020, both stocks generally increased, displaying a notably stronger correlation (evident from the medium green scatters’ slope in Figure 1). This change is possibly linked to Pfizer’s vaccine development positively impacting travel confidence, thereby affecting the aviation sector. In 2022, this correlation moderated, aligning with the evolving COVID-19 situation. By 2023, as COVID-19 concerns eased, the correlation between PFE and BA stock prices significantly weakened, approaching a negligible level (indicated by the lightest yellow scatters in Figure 1).
Conducting a linear regression on the data from different years (as shown by the varying scatter colors in Figure 1) reveals distinct slopes and intercepts for each year. The slope, though not a precise measure, suggests the causal influence of PFE on BA, while the intercept reflects BA’s price shifts attributable to other factors. These variations highlight a clear nonstationary relationship between the two companies.
@@ 3.2 From Causal Discovery Perspective
4 Modeling and Discovery of Causally Related Latent Factors
4.1 From Financial Perspective
@@
In stock market prediction, it has been shown by Fama and French Fama and French (1996) that, their classical 3-factor model, which consists of the market factor, a factor based on market equity (small-minus-big, SMB), and a factor based on bookto-market equity (high-minus-low, HML), well summarizes the cross section of average stock returns as of the mid-1990s. It works well empirically Hou et al. (2017) though it assumes linearity relations among variables. However, it has become clearer recently that the Fama-French model fails to account for a wide array of asset pricing anomalies Hou et al. (2015).
Therefore, over the past 2 decades, researchers have been devoted to refining the 3-factor model, with prominent ones like the 4-factor model Hou et al. (2015) and the 5-factor model Fama and French (2015). This line of thought can be summarized as finding the variables that are essential for prediction but currently unknown or cannot be observed, and then incorporating them for prediction. For example, the 3-factor model Fama and French (1996) predicts the return of the $i$-th asset by $r_{t}^{i}-r_{t}^{f}=\alpha_{t}^{i}+\beta_{MRP}^{i}MRP_{t}+\beta_{SMB}^{i}SMB_{t}+\beta_{HML}^{i}HML_{t}+\epsilon_{t}^{i}$, where MRP is the market risk premium, SMB is the size premium, and HML is the value premium; based on that, the five-factor model incorporates two factors RMW, the difference between the returns on diversified portfolios with robust and weak profitability, and CMA, the difference between the returns on diversified portfolios of low and high investment firms.
Following this line of thought, it can be argued that one promising direction toward better prediction in the area would be unveiling the factors that are essential in the underlying dynamics but cannot be explained by the existing factors found by the field. We refer to these variables/factors as latent variables. It is possible that part of the latent variables relate to some existing concepts and can be measured, but we have no observation of them just because we do not know what are they. At the same time, some latent variables might relate to some new concepts in the field. We argue that, with cutting-edge causal discovery methods at hand, we are able to find the existence and cardinality of these latent variables, and even the causal relations among all the variables.
4.2 From Causal Discovery Perspective
In this paper, we delve into the fundamental task of analyzing financial data: identifying causal relationships among different stocks to determine influence patterns. This process is known as causal discovery, i.e., to recover the causal structure that delineates the causal relations among variables from observed data.
Causal discovery typically identifies structure by leveraging (conditional) independence relations. Rooted in Reichenbach’s Common Cause Principle (Hofer-Szabó et al., 1999), it posits that two variables are interdependent due to either a direct causal relation or a shared common cause, which if observed, will render them conditionally independent. Many early causal discovery methods, such as PC (Spirtes et al., 2000) and GES (Chickering, 2002), thus operate under the causal sufficiency assumption, i.e., assuming the absence of unmeasured common causes among observed variables, enabling accurate identification of the causal structure (up to an equivalence class).
However, real-world data often violate the causal sufficiency assumption, exemplified by the factor model discussed above in financial data, and also by viewing the changing quantitie that drives nonstationarity as latent variables. The presence of unmeasured confounders can lead to spurious causations when applying methods like PC and GES directly on observed data. These dependencies among variables remain unaccounted for by their confounders. To address this, approaches like Fast Causal Inference (FCI (Spirtes et al., 2000)) and its variants (Spirtes et al., 2013; Colombo et al., 2012; Akbari et al., 2021) are developed to leverage conditional independence tests to allow for causal discovery in the presence of unmeasured confounders.
Nevertheless, FCI’s limitation lies in that its results on observed variables may be too generalized and lack meaningful interpretation due to large indeterminacies (e.g., whenever it is possible for two measured variables to be confounded, it will indicate so). Moreover, it focuses solely on the causal relations among observed variables, without considering the relations among latent variables. It only allows the existence of unobserved confounders but has to assume their mutual independence. This limitation is also shared by methods using over-complete ICA-based techniques like (Hoyer et al., 2008; Salehkaleybar et al., 2020). Then, what if there are also causal relations among latent variables, where our interest also extends to? The following example provides an introduction to addressing this challenge:
Example 2
Suppose we observe four variables ${X_{1},X_{2},X_{3},X_{4}}$ as common children of a latent variable $L$. They adhere to a linear structural equation model (SEM) as follows:
[…]
$X_{1}$ $=aL+E_{1}$
$X_{2}$ $=bL+E_{2}$
$X_{3}$ $=cL+E_{3}$
$X_{4}$ $=dL+E_{4}$
where ${E_{1},E_{2},E_{3},E_{4}}$ are mutually exogenous noise components for each variable. In the absence of observable $L$, the observed variables ${X_{1},X_{2},X_{3},X_{4}}$ show no (conditional) independencies Applying PC, GES, or FCI directly will thus result in a fully connected graph, where there is no (conditional) independencies either. Then, is it possible to distinguish between the two cases by utilizing some statistical information other than (conditional) independencies? Fortunately, the non-adjacency among observed variables leaves trace as rank deficiency in their variance-covariance matrix: e.g., consider the variance-covariance matrix between $\{X_{1},X_{2}\}$ and $\{X_{3},X_{4}\}$, $\begin{bmatrix}ac&ad\\ bc&bd\end{bmatrix}\cdot\operatorname{cov}(L)$ – it is a singular matrix. In contrast, a truly fully connected graph would not exhibit such rank deficiency.
This example illustrates the so-called Tetrad constraint, exploits the assumption of linear causal relations in the structural equation model to identify latent variables. The rank-deficiency among the observed variables often indicate the existence of latent variables, which, can be understood as another way render the (conditional) independencies among observed variables, though the conditioned one is not available. Following Tetrad constraint, methods have been developed (Silva et al., 2006; Silva and Scheines, 2005; Sullivant et al., 2010; Spirtes, 2013b; Kummerfeld and Ramsey, 2016; Huang et al., 2022), albeit often under strict structural assumptions (e.g., at least two observed children for each latent variable). To mitigate these constraints, alternative methods have been developed to utilize e.g., higher-order information (non-Gaussianity) (Shimizu et al., 2009; Cai et al., 2019; Xie et al., 2020; Adams et al., 2021; Dai et al., 2022; Chen et al., 2022), matrix-decomposition (Anandkumar et al., 2013), and mixture oracles (Kivva et al., 2021). However, as a trade-off to better identification results, these methods require more parametric assumptions on the underlying causal process. In this paper, parametrically we only assume the linearity of causal relationships, and investigate what information can be acquired with only rank deficiency constraints.
@@ 4.3 Causal Discovery in the Presence of Latent Variables
Latent variables are ubiquitous and essential in machine learning. For example, pixels of an image are dependent and yet it can be argued that they are confounded by the latent underlying concept. Other examples would be causally related concepts in psychological studies: we are interested in the human hidden mental factors and yet we can only rely on questionnaires to measure the latent mental states. Generally speaking, identifying causal relations among latent variables is hard. Traditional causal discovery methods such as PC (Spirtes et al., 2000) have to assume the absence of latent confounders. The FCI (Spirtes et al., 2000) asymptotically returns the correct causal relations among measured variables, even in the presence of latent ones, but it is generally not informative: it does not try to locate causally related latent variables and cannot identify the relations that involve latent variables. Recent advances in overcomplete ICA provide a way to handle the latent causal discovery problem, but it assumes latent variables are independent, and doesn’t provide unique solutions in general.
Therefore, there is an urgent need for a computational method that can learn true causality including causally related latent variables and causal relations. Although it is well known to be an extremely hard problem, interestingly, our recent preliminary results have shown it can be reliably achieved by exploring certain distributional properties of measured variables - by properly making use of rank deficiency and GIN constraints. Specifically one can recover latent variables and involve the causal relations by making use of rank deficiency of covariance matrix (Dong et al., 2023) or testing for independence between combinations of measured variables known as GIN (Xie et al., 2020). We provide our solution as follows (more details can be found in our manuscript (Dong et al., 2023)).
@@
We further have the rank faithfulness assumption for latent linear causal models as follows. A probability distribution $p$ is rank faithful to $\mathcal{G}$ if every rank constraint on a sub-covariance matrix that holds in $p$ is entailed by every linear structural model with respect to $\mathcal{G}$. It is the classical faithfulness assumption that is critical and prevalent in causal discovery (Spirtes et al., 2000; Huang et al., 2022; Dong et al., 2023); it holds generically on infinite data, as the set of values of the SCM’s free parameters for which rank is not faithful is of Lebesgue measure 0 (Spirtes, 2013a).
Goal. Given i.i.d. samples of observed variables $\mathbf{X}_{\mathcal{G}}:=\{\mathsf{X_{i}}\}_{i=1}^{n}$ generated per LLCM, we want to (i) locate all the latent variables $\mathbf{L}_{\mathcal{G}}$, determine their cardinalities, and identify the causal structure $\mathcal{G}$ over all the variables, including both latent and observed ones. To fulfill (i), we will leverage the information from the rank of the covariance matrix of observed variables.
Rank Constraint and Motivation. We propose to leverage information from the rank of the cross-covariance matrix over observed variables to unveil the underlying causal structure. The intuition lies in that t-separations are good graphical indicators and they can be inferred from rank constraints. We shall first introduce the definition of trek and t-separation, why d-separation is a special case of t-separation, as well as how t-separations are related to rank constraints. Then we will discuss why t-separations are particularly useful in the presence of latent variables.
@@
Now we show why t-separation implied by rank is a good graphical indicator, especially in the presence of latent variables. Take the graph in Example 2 as an example. If we can observe $\mathsf{L}$ then we can directly do the CI test conditional on $\mathsf{L}$, and thus find all the structure. However, checking that CI relation is impossible, and by CI tests we can only find a clique among $\mathsf{X_{1}},\mathsf{X_{2}},\mathsf{X_{3}},\mathsf{X_{4}}$. Fortunately, t-separation implied by rank constraints can achieve a similar goal when $\mathsf{L}$ is not observed. Specifically, we can observe that $\text{rank}(\Sigma_{\{\mathsf{X_{1}},\mathsf{X_{2}}\},\{\mathsf{X_{3}},\mathsf{X_{4}}\}})=1$, which implies the existence of one variable being in the treks between $\{\mathsf{X_{1}},\mathsf{X_{2}}\}$ and $\{\mathsf{X_{3}},\mathsf{X_{4}}\}$. The rationale behind is that the t-separation of $\mathbf{A}$, $\mathbf{B}$ by $(\mathbf{C}_{\mathbf{A}},\mathbf{C}_{\mathbf{B}})$ can be deduced through rank without observing any element in $(\mathbf{C}_{\mathbf{A}},\mathbf{C}_{\mathbf{B}})$. This property makes t-separations particularly useful for unveiling the underlying causal structure when latent variables exist. Next, we formalize this intuition into the concept of atomic cover (a cover is a set of variables) as the minimal identifiable substructure in a graph, with effective cardinality of a set of covers $\mathcal{V}$ defined as $||\mathcal{V}||=|(\cup_{\mathbf{V}\in\mathcal{V}}\mathbf{V})|$.
@@
    There does not exist a partition of $\mathbf{V}=\mathbf{V_{1}}\cup\mathbf{V_{2}}$ such that both $\mathbf{V_{1}}$ and $\mathbf{V_{2}}$ are atomic covers.
The atomic cover is essential as it is the minimal identifiable substructure in a graph. Specifically, given the following Condition 6, we can strive to ensure that the rank deficiency property uniquely implies an atomic cover, and this is formally captured by Theorem 7 as follows.

### 12 Jan 2024 v2

@@ 1 Introduction and Related Work
Another focal point in causal discovery and finance is the issue of data nonstationarity, influencing the outcomes of causal discovery methods applied to financial data. Nonstationarity implies that causal relations and their strengths vary with time (Zhang et al., 2015). Within finance, this phenomenon is commonly referred to as concept drift, signifying the changing distribution of data over time (Lu et al., 2018). Traditional causal discovery methods such as PC (Spirtes et al., 2000), GES (Chickering, 2002), and LiNGAM (Shimizu et al., 2006) were originally designed for the i.i.d. case and did not account for distribution shifts. A recent method, PC-MCI (Runge et al., 2019), is specifically tailored for time series data; however, it still exhibits limitations when confronted with changing causal relations over time.
@@
Our contributions can be summarised as below:
  • We delineate three pivotal issues concerning causality in financial data, i.e., time resolution, nonstationarity, and latent variables, and undertake a comprehensive exploration on these challenges following a systematic and causal point of view.
  • We provide a novel and inspiring causal understanding of the three issues. Based on this understanding we propose rigorous methods to address each issue as well as a systematic solution that works even when the three issues are present simultaneously.
  • Our empirical study validates the effectiveness of our proposed methods in real-life scenarios in finite sample cases. Our techniques and findings hopefully shed new light on the problem and serve as a foundation for further exploration in the area.
@@ 1.1 Concepts and Terminology in Causal Discovery
Structural Causal Model. The Structural Causal Model (SCM) (Pearl, 2009) comprises a set of endogenous variables ($V$) and a set of exogenous variables ($U$), interconnected by a set of functions ($F$) that illustrate how the changes of exogenous variables $U$ impact the values of endogenous variables $V$ in the data generation process. For instance, for $x_{i}\in V$, $u_{i}\in U$, and $f_{i}\in F$, the model $x_{i}:=f_{i}(Pa(x_{i}),u_{i}),i=1,\dots,d$, signifies the assignment of the value $x_{i}$ to a function $f_{i}$ of its structural parents $Pa(x_{i})$ and exogenous variable $u_{i}$. For each SCM, a causal graph $\mathcal{G}$ is correspondingly generated by representing each $x_{i}$ with each vertex and adding directing edges from each parent variable in $Pa(x_{i})$ to the child $x_{i}$.
Causal Sufficiency. A set of observed variables is deemed causally sufficient if all common causes of these variables are also observed (Spirtes et al., 2000). In other words, there is no latent confounder in the graph and all causally related variables are observed.
Causal Markov Condition. In the causal graph, each variable is independent of its non-descendants conditional on all of its direct causes (Hausman and Woodward, 1999). This relationship establishes a connection from the causal graph to the statistical properties.
Markov Equivalence Class. A set of directed acyclic graphs belong to the same Markov Equivalent Class if they entail the same conditional independence relations among the observed variables (Pearl et al., 2000; Spirtes et al., 2000). Elements in the same class have the same causal skeleton and v-structures.
Faithfulness. A probability distribution P is faithful to a causal graph $\mathcal{G}$ if all conditional independence relations in the probability distribution P are entailed by the causal graph $\mathcal{G}$ (Spirtes et al., 2000). This relationship establishes a connection from the statistical properties to the causal graph.
@@ 2 Time Resolution and Aggregation
$\mathbf{X}_{t}=\mathrm{A}\mathbf{X}_{t-1}+\mathbf{e}_{t},$ (1)
where $\mathbf{X}_{t}=(X_{t,1},\dots,X_{t,n})^{T}$ represents the original data vector, $\mathbf{e}_{t}=(e_{t,1},\dots,e_{t,n})^{T}$ is the temporally and contemporaneously independent noise vector, and $\mathrm{A}$ is the causal transition matrix. Furthermore, the temporally aggregated data (observed data)
@@
Taking Eq. 1 and Eq. 2 together, we have:
$\tilde{\mathbf{X}}_{t}$ $=\frac{1}{k}\sum_{i=1}^{k}\mathbf{X}_{i+(t-1)k}=\frac{1}{k}\sum_{i=1}^{k}\left(\mathrm{A}\mathbf{X}_{i+(t-1)k-1}+\mathbf{e}_{i+(t-1)k}\right)$ (3)
$=\mathrm{A}\left[\frac{1}{k}\sum_{i=1}^{k}\mathbf{X}_{i+(t-1)k}-\frac{1}{k}\left(\mathbf{X}_{tk}-\mathbf{X}_{(t-1)k}\right)\right]+\frac{1}{k}\sum_{i=1}^{k}\mathbf{e}_{i+(t-1)k}.$
@@
$\tilde{\mathbf{X}}_{t}=\mathrm{A}\tilde{\mathbf{X}}_{t}+\tilde{\mathbf{e}}_{t},$ (4)
where $\tilde{\mathbf{e}}_{t}=\frac{1}{k}\sum_{i=1}^{k}\mathbf{e}_{i+(t-1)k}$ is the mixture of independent component vectors $\mathbf{e}_{t}$, and elements in $\tilde{\mathbf{e}}_{t}$ are contemporaneous independent to each other. Clearly, $\tilde{\mathbf{X}}_{t}$ exhibits behaviors resembling a linear instantaneous causal process in the i.i.d. case, and hence instantaneous causal discovery methods can be directly applied to aggregated data to discover meaningful causal relationships (detailed in Section 4).
@@ 3.1 From the Financial Perspective
Figures 1 demonstrates the evolving correlation between the stock price daily log return time series of Pfizer Inc. (PFE) and The Boeing Company (BA) from 2019 to 2023. This relationship has experienced significant shifts, particularly during the pre- and post-COVID periods. In 2019, a time of economic expansion, both PFE and BA exhibited an upward trend in stock prices. The correlation observed in this era could be influenced by broader economic factors (confounders). The correlation is relatively weak in this period. The onset of COVID-19 in early 2020 saw a sharp decline in both BA’s and PFE’s stock prices. However, in late 2020, both stocks generally increased, displaying a notably stronger correlation (also demonstrated by the correlation between the log daily returns in 2020). This change is possibly linked to Pfizer’s vaccine development positively impacting travel confidence, thereby affecting the aviation sector. After 2021, this correlation moderated, aligning with the evolving COVID-19 situation. By 2023, as COVID-19 concerns eased, the correlation between PFE and BA stock prices significantly weakened.
Of course, it’s worth noting that pairwise relationships don’t always reflect true causality (may be subject to confounders), which is typically captured by conditional relationships. Therefore, the explanations provided above serve more for intuitive understanding than accurate discovery. For a more systematic investigation of this example, please refer to Section 5.
Figure 1: Example to illustrate the nonstationary relationship between Pfizer Inc. (PFE) and The Boeing Company (BA). The upper row: five scatter plots correspond to years from 2019 to 2023. The scatters are the log daily returns of the two stocks. Clearly they have different dependence patterns. The linear regression line and the correlation are annotated. The lower row: daily prices series of the two stocks are plotted for reference.
@@ 3.2 From Causal Discovery Perspective
4 Finding Causal Model in the Presence of Latent Factors
4.1 Factor Model from A Causal Perspective
@@
In stock market prediction, it has been shown by Fama and French Fama and French (1996) that, their classical 3-factor model well summarizes the cross section of average stock returns as of the mid-1990s. Over the past 2 decades, researchers have been devoted to refining the 3-factor model, with prominent ones like the 4-factor model Hou et al. (2015) and the 5-factor model Fama and French (2015). Though this line of research works well in some scenarios (Hou et al., 2017), it has become clearer recently that it fails to account for a wide array of asset pricing anomalies (Hou et al., 2015).
One can continuously add more factors to enhance the model, and yet, we argue that, without a fundamental causal understanding of the problem, one can never make reliable and interpretable predictions based on a factor model. The reason lies in that the real objective of a factor model, say, $\mathsf{Y}=a\mathsf{X}+b\mathsf{Z}+\epsilon$, is to predict the causal effect $\mathbb{E}[\mathsf{Y}|\text{do}(\mathsf{X})]$, rather than the conditional expectation $\mathbb{E}[\mathsf{Y}|\mathsf{X},\mathsf{Z})]$ (Lopez de Prado, 2022). Therefore, without a causal perspective, one cannot even correctly formulate the quantity that we really care about. Even though we might correctly formulate the problem, it is still very hard, if not impossible, to predict the causal effect, as generally we only have access to observational data.
Fortunately, with advanced causal discovery methods at hand, one can identify the causal graph and then perform simulated interventions (e.g., backdoor adjustment(Pearl, 2009)) to estimate the causal effect, by making use of only observational data. Typical causal discovery methods include the constraint-based PC algorithm (Spirtes et al., 2000) and score-based GES (Chickering, 2002), both of which can asymptotically arrive at the Markov equivalence class of the underlying causal graph, by assuming causal sufficiency.
However, latent variables are ubiquitous in finance data, and thus causal sufficiency generally fails to hold in the area. Roughly speaking, latent variables/factors are those variables that are essential in the underlying causal dynamics, but cannot be directly observed, or even cannot be semantically well captured by existing concepts in the field. For example, policies taken by governments can have a significant impact on a financial system, and yet some policies may not always be directly observable. Therefore, causal discovery methods that work in the presence of latent variables would be very favorable. One famous line of thought that allows latent variables is FCI (Spirtes et al., 2000) and its variants (Spirtes et al., 2013; Colombo et al., 2012; Akbari et al., 2021). It makes full use of CI information and correctly identifies causal relations between observed variables.
Although FCI provides correct graphical information in the presence of latent variables, its results tend to be too generalized due to large indeterminacies. E.g., whenever it is possible for two measured variables to be confounded, it will indicate so. Moreover, it focuses solely on the causal relations among observed variables, without considering the relations among latent variables. Recent advances in overcomplete ICA (Hoyer et al., 2008; Salehkaleybar et al., 2020) also aim at solving this problem, but they have to assume all latent variables are independent and they do not provide unique solutions in general.
Therefore, a question naturally arises. If we allow all the variables including observed and latent ones to be very flexibly related (e.g., a latent variable can serve as the effect, the cause, or the mediation of any variables), can we still identify the whole underlying causal structure (including the location, cardinality, and relations among all the variables), and thus gain a better understanding of the underlying causal model? The answer is yes and we shall detail it in the following section.
@@ 4.2 Causal Discovery in the Presence of Latent Variables
Latent variables are ubiquitous and essential in many fields. For example, pixels of an image are dependent and yet it can be argued that they are confounded by the latent underlying concept. Other examples would be causally related concepts in psychological studies: we are interested in the human hidden mental factors and yet we can only rely on questionnaires to measure the latent mental states. Therefore, there is an urgent need for a computational method that can learn true causality including causally related latent variables and causal relations.
Although it is well known to be an extremely hard problem, interestingly, our recent preliminary results have shown it can be reliably achieved by exploring certain distributional properties of measured variables - by properly making use of rank deficiency and generalized independence noise (GIN) constraints. Specifically one can recover latent variables and involve the causal relations by making use of rank deficiency of covariance matrix (Dong et al., 2023) or testing for independence between combinations of measured variables known as GIN (Xie et al., 2020). We provide our solution (more details can be found in our manuscript (Dong et al., 2023)), and start with the definition of latent linear causal model as follows.
@@
Goal. In the above definition, we have $n+m$ variables in the causal model and only n out of n+m are observed variables. Given i.i.d. samples of only observed variables $\mathbf{X}_{\mathcal{G}}:=\{\mathsf{X_{i}}\}_{i=1}^{n}$, our goal is to locate all the latent variables $\mathbf{L}_{\mathcal{G}}=\{{L}_{i}\}_{i=1}^{m}$, determine their cardinalities, and identify the causal structure $\mathcal{G}$ over all the variables, including both latent and observed ones. To this end, we will leverage the information from the rank of the covariance matrix of observed variables and will assume rank faithfulness for latent linear causal models as follows. A probability distribution $p$ is rank faithful to $\mathcal{G}$ if every rank constraint on a sub-covariance matrix that holds in $p$ is entailed by every linear structural model with respect to $\mathcal{G}$. It is the classical faithfulness assumption that is critical and prevalent in causal discovery (Spirtes et al., 2000; Huang et al., 2022; Dong et al., 2023); it holds generically on infinite data, as the set of values of the SCM’s free parameters for which rank is not faithful is of Lebesgue measure 0 (Spirtes, 2013).
Rank Constraint and Motivation. We propose to leverage information from the rank of the cross-covariance matrix over observed variables to unveil the underlying causal structure. The intuition lies in that t-separations are good graphical indicators and they can be inferred from rank constraints. We begin with a motivating example as follows.
Example 2
Suppose we observe four variables $\mathsf{X_{1}},\mathsf{X_{2}},\mathsf{X_{3}},\mathsf{X_{4}}$ as common children of a latent variable $\mathsf{L_{1}}$. They adhere to a LLCM (Definition 1) as follows:
$\mathsf{X_{1}}$$\mathsf{X_{2}}$$\mathsf{X_{3}}$$\mathsf{X_{4}}$$\mathsf{L_{1}}$$a$$b$$c$$d$
$\mathsf{X_{1}}$ $=a\mathsf{L_{1}}+\epsilon_{1}$
$\mathsf{X_{2}}$ $=b\mathsf{L_{1}}+\epsilon_{2}$
$\mathsf{X_{3}}$ $=c\mathsf{L_{1}}+\epsilon_{3}$
$\mathsf{X_{4}}$ $=d\mathsf{L_{1}}+\epsilon_{4}$
The classical line of thought for causal discovery is to use CI tests to find d-separations in the graph. E.g., if we assume that $\mathsf{L_{1}}$ is observed, we can simply get $\mathsf{X_{1}}\perp\!\!\!\perp\mathsf{X_{2}}|\mathsf{L_{1}}$ and thus conclude the skeleton among these three variables. However, $\mathsf{L_{1}}$ is hidden, and thus we cannot find any CI relations from observational data. by using CI information we can only arrive at a fully connected graph among $\mathsf{X_{1}},\mathsf{X_{2}},\mathsf{X_{3}},\mathsf{X_{4}}$.
By the above example, we show that in the presence of latent variables, CI relations only provide limited information about the underlying graph. Next, we will show that by using t-separation implied by rank constraint, we are able to identify the latent variable, which is beyond what CI can do. We shall begin with the definition of trek and t-separation.
@@
Readers who are less familiar with t-separation may refer to our manuscript (Dong et al., 2023) for illustrative examples. Just as d-separation can be implied by CI, t-separation can be implied by rank of the cross-covariance matrix over observed variables, as follows.
@@
Now we show why t-separation implied by rank is a good graphical indicator, especially in the presence of latent variables. Again consider the graph in Example 2. If we can observe $\mathsf{L_{1}}$ then we can directly do the CI test conditional on $\mathsf{L_{1}}$, and thus find the skeleton. However, checking that CI relation is impossible, and by CI tests we can only find a clique among $\mathsf{X_{1}},\mathsf{X_{2}},\mathsf{X_{3}},\mathsf{X_{4}}$. Fortunately, t-separation implied by rank constraints can achieve a similar goal when $\mathsf{L}$ is not observed. Specifically, we can observe that $\text{rank}(\Sigma_{\{\mathsf{X_{1}},\mathsf{X_{2}}\},\{\mathsf{X_{3}},\mathsf{X_{4}}\}})=1$, which implies the existence of one variable being in the treks between $\{\mathsf{X_{1}},\mathsf{X_{2}}\}$ and $\{\mathsf{X_{3}},\mathsf{X_{4}}\}$. The rationale behind is that the t-separation of $\mathbf{A}$, $\mathbf{B}$ by $(\mathbf{C}_{\mathbf{A}},\mathbf{C}_{\mathbf{B}})$ can be deduced through rank without observing any element in $(\mathbf{C}_{\mathbf{A}},\mathbf{C}_{\mathbf{B}})$.
On the other hand, if we cannot observe $\mathsf{X_{4}}$, we cannot identify the existence of latent variable $\mathsf{L_{1}}$. The reason lies in that if we cannot observe $\mathsf{X_{4}}$, we cannot calculate $\text{rank}(\Sigma_{\{\mathsf{X_{1}},\mathsf{X_{2}}\},\{\mathsf{X_{3}},\mathsf{X_{4}}\}})$. Instead, we can only test $\text{rank}(\Sigma_{\{\mathsf{X_{1}},\mathsf{X_{2}}\},\{\mathsf{X_{3}}\}})$ and it will always be 1, no matter whether there is a latent variable or not. Therefore, we have the rough intuition that, for a latent variable to be identifiable, it must have enough observed children and neighbors. Next, we formalize this intuition into the concept of atomic cover (a cover is a set of variables) as the minimal identifiable substructure in a graph, with effective cardinality of a set of covers $\mathcal{V}$ defined as $||\mathcal{V}||=|(\cup_{\mathbf{V}\in\mathcal{V}}\mathbf{V})|$, and that $\mathbf{Y}$ are pure children of $\mathbf{X}$, i.e., $\mathbf{Y}\in\text{PCh}_{\mathcal{G}}(\mathbf{X})$, iff $\text{Pa}_{\mathcal{G}}(\mathbf{Y})=\cup_{\mathsf{Y_{i}}\in\mathbf{Y}}\text{Pa}_{\mathcal{G}}(\mathsf{Y_{i}})=\mathbf{X}$ and $\mathbf{X}\cap\mathbf{Y}=\varnothing$.
@@
    There is not a partition of $\mathbf{V}=\mathbf{V_{1}}\cup\mathbf{V_{2}}$, s.t., both $\mathbf{V_{1}}$ and $\mathbf{V_{2}}$ are atomic covers.
The concept of atomic cover is essential as it is the minimal identifiable substructure in a graph. Specifically, given the following Condition 6, we have that the rank deficiency property uniquely implies an atomic cover, formally captured by Theorem 7.